Thats it!! i can get c6 colombina! i just have to be a little lucky
Commentaires
Raymond490902026-01-13 18:34
Go out into a thunderstorm to farm some RNG. Every time you don't get struck by lightning means the odds will even out eventually, right?
MiniMages2026-01-13 18:46
But if they dodge 6 lightning strikes they'll have used up their luck.
Puzzlehead_Lemon2026-01-13 19:06
Nah, they have to dodge 100 in a row to get something to unlock their ultimate weapon. (Yes, a Final Fantasy X reference in the wild!)
MiniMages2026-01-13 19:13
Hahahaha, it's not as bad as trying to calibrate the towers in X-2.
Puzzlehead_Lemon2026-01-13 21:55
I blocked that out of my memory completely. I'm going to go stare off into the middle distance for a while and listen to Fortunate Son
jasonensteinyt2026-01-13 18:51
That's if he doesnt get *torn into oblivion*
thwrlsgenshin2026-01-13 19:02
beware with *rats from the sky*
Trebord_2026-01-13 19:57
#INAZUMA SHINES ETERNAL
intadtraptor2026-01-13 20:43
Final Fantasy X Lulu Weapon Farming Flashbacks.
ImNotTiredOfWinning2026-01-13 18:36
With those odds, you can go play the powerball, win, and then get just enough money to top-up and pull for her C6
Low_Artist_76632026-01-13 18:41
"just do real gambling to pay for fake one"
Nexius_2026-01-13 19:05
they don't want you to know this simple trick
ImNotTiredOfWinning2026-01-13 20:32
You saw that Pop-Up ad Too?
SaibaShogun2026-01-14 00:23
“But that’s how losers think.”
DreamweaverMirar2026-01-13 20:05
Genius
arminismyspirtanimal2026-01-13 18:39
Perhaps your the MC of a slice of life anime who is just incredibly lucky that it defies all odds
Puzzlehead_Lemon2026-01-13 19:09
Just watch out for Truck-kun. "I C6'd Columbina in a single pull, then a truck crashed into my house and killed me, now I'm grinding goblins in another world to survive... but in a strip club?!"
WorstSkilledPlayer2026-01-13 20:58
😂😂😂. Truck-kun remain undefeated.
WastePermission96202026-01-14 00:16
“I maxed out a character in a single pull and now I must romance her in her home world?!?”
Darkfire13782026-01-13 18:40
Lotto got better odds.
TinderGirl922026-01-13 18:41
if you go for c6 columbia then you also need R5 signature weapon. otherwise its just a c6 columbina
SoC1752026-01-13 18:43
Isn't it even worse as you need at least 7 featured copies? Pulling a "mere 6" only gets you C5
kissbence992026-01-13 18:49
Thats wrong, you need at least 7 copies, so the chance is \~0.000000000000026%
app082026-01-13 20:04
Okay but that's still a chance
Hougasej2026-01-13 20:52
Thats for exact 7 only, for "at least 7" or 7+ copies chance is 0,0000000000033%
P(X>=7) = P(7) + P(8) + P(9) + P(10)
Edit: Wait a second, i calculated wrong! This (3.3×10^-14) astronomically low chance is for any 5 stars, not for limited. I didnt include 50/50 and guarantee. For c6 limited char chance is 0,000000000000000257%, or 2.57×10^-16
kissbence992026-01-13 21:06
Mine is correct, I think you ignored 5050
Hougasej2026-01-13 21:33
Yeah, i totally forgot about it, but your result is too high, its not correct. Your result is too near for Bernoulli P = 10C7 × 0.006^7 × 0.994^3, but pity system making it like 128 times worse so right chance is lower, 2.57×10^-16
5mil_2026-01-13 21:51
account for the fact that if one of the 3 other ones is a 50/50 loss, then the next 5 star is guaranteed
(significantly more effort but whatever)
oh and also capturing radiance which lowers the chance of 3 consecutive 50-50 losses, but the effect of any of these is minimal
your 0.994^3 or 0.997^3 assumes that the others can't be 5 stars, which they can, so it should just be 1^3 which is 1 (oh right you did say it earlier)
Hougasej2026-01-14 06:24
Yeah, im again forgot something, this time P(X>=7) = P(7) + P(8) + P(9) + P(10) i wrote myself earlier, lol. P(X>=7) = 3.3×10\^-14 + 7.5×10\^-17 + 1.0×10\^-19 + 6.0×10\^-23 = 3.307510006 × 10\^-14
P(7) chances to c6 is 0.78%(1/128 i wrote earlier, need 7 wins), P(8) chances to c6 is 6.25%(1/16 cause 2×0.5\^8 of 8 wins and 7 wins and 1 lose at the end, and 7 variants of 0.5\^7 with guarantee in the middle, with sum of 1/256+1/256+7/128=16/256=1/16 or 6.25%), p(9) and p(10) contribution is negligible so i just ignored it. Just sum it and real final chance to c6 in 10 pulls is (3.3×10\^-14)×0.78% + (7.5×10\^-17)×6.25% = 2.625×10\^-16
edit: recalculated, firstly got wrong result. Btw, why this time formatting doesn't work??? Btw my brain is fried, i edited it like 10 ten time after each recalculation.
kissbence992026-01-13 22:04
First, my previous reply sounds passive aggressive, so I want to apologize for that :D
[This](https://docs.google.com/spreadsheets/d/1XXbShUXR-RLmDxTZmXzl_Z7ye5ve1lfaVmLHXsKFvxg/edit?usp=sharing) was my calculation, let me know if I did this wrong. Technically my calculation doesn't take into account the probability of getting a guaranteed after losing 5050 on one of the 3 pulls that we can spare, but I don't think that changes the results in a meaningful way.
I'm not sure how the pity system comes into play, we are always before any kind of soft pity, so every pull should have a chance of 0.3% for the 5 star we want.
Hougasej2026-01-14 06:54
i just used Bernoulli formula for mine calculations. Pity is kinda easy to calculate for this exact scenario, cause highest possible scenario P(7) requires 7 5050 wins for 7 tries, second highest P(8) let you lose 1 time(so it either 8 wins, 7 wins and 1 lose at the end, or guarantee at the middle, so 1/256+1/256+7/128=16/256=1/16 or 6.25%), and P(9) and P(10) is so low that you can basically ignore them. I mean, you can calculate P(7), then multiply it to 0.5\^7(0.0078125 or 0.78125%, or just divide to 128, it same thing), then calculate P(8), multiply to 6.25% from earlier calcs, sum them - and you got your right final chance.
I see you used binomial distribution which is also right thing, but may underestimate chances in compare to bernoulli(not this time). You just need to use right chances and results become similar with mine. (i made copy of your doc, changed chance from 0.3% to 0.6%, included pity and got 2.625e-16)
edit: forget what i wrote in edit earlier, right calculations is above
Hougasej2026-01-14 07:42
I feel like i may be wrong on smth again so asked gemini pro to recalculate using python, it got something near my calcs:
Code execution result
Five Stars | Binomial Prob | Chance of 7+ Event chars | Contribution
--------------------------------------------------------------------------------
7 | 3.29913e-14 | 0.00781 | 2.57744e-16
8 | 7.46784e-17 | 0.06250 | 4.66740e-18
9 | 1.00172e-19 | 0.22656 | 2.26953e-20
10 | 6.04662e-23 | 0.50000 | 3.02331e-23
--------------------------------------------------------------------------------
Total Probability: 2.6243449537348796e-16
One in 3,810,474,680,841,151
Kaymish_2026-01-13 21:59
So you're telling me there's a chance?
WandererMisha2026-01-13 23:08
9 copies to get her to lvl 100!
beautheschmo2026-01-14 02:28
But it isn't zero, and for me that makes it the same as 100%
DarkAlatreon2026-01-13 18:50
AI content, cool
Similar_Medicine_8642026-01-13 19:12
I stg people these days do not understand the importance of ai. They instantly assume ai=bad.
brliron2026-01-13 19:26
Well, in this specific case, the AI made a mistake, and OP didn't notice. So, yeah, ai=bad.
Ecchify2026-01-13 19:42
mh yes let's waste natural resources, promote not using our brains, increase the ram prices by a billion percent....all for it to not even give you the right answer
the number it gave is wrong, literally just input that on a calculator to check. and the last statement about lightning odds don't make any sense either lmao
cmon man, let's use our brains a little. literally no one besides billionaires benefit from dumb mindsets like yours
Velocity1112026-01-13 20:24
I'm not saying the AI did it right, but it's not so easy to just put this calculation in a calculator... in order to calculate the odds of c6 columbina in a 10 pull, you have to add up the odds of 7/10 5 stars all winning the 50/50, 8/10 5 stars with 1 50/50 loss and with 0 50/50 loss, 9/10 5 stars with 2, 1, and 0 50/50 loss, and 10/10 5 stars with 3, 2, 1, and 0 50/50 loss, taking into consideration the radiant capture possibilities as well... it's a lot of calculation
Ecchify2026-01-13 20:33
dont really get your point here?
and 10 pull is an arbitrary number anyway, if you want a simple calculation that doesn't need any knowledge about maths just calculate odds for 7 in a row instead.
Velocity1112026-01-13 20:39
the point is that "just put it in a calculator" is not so straightforward for the question the OP was trying to answer. The odds of 7 5 stars in a row is not the same as the odds of at least 7 on-banner 5 stars in a 10 pull, and the real answer could be larger or smaller by a couple orders of magnitude.
ultimately it doesn't matter either way since the answer is some super small number that in practical terms is just 0, but if someone actually wanted to know the real answer to that question for some reason, it's not a straightforward calculation
edit: at least 7, not exactly 7
Ecchify2026-01-13 20:49
i'll be frank, if you're that curious about something that you can figure out with 5 minutes of research, you should probably do that. and if you're too lazy then just ask, someone might be interested in finding that out themselves. instead of using shit that's harmful in many ways
Velocity1112026-01-13 20:52
the first thing I said was "I'm not saying the AI is right", so yeah I'm not saying AI is the correct thing to use here. The only thing I'm saying is you're dismissing this as an easy thing to calculate or to plug into a calculator that, in all likelihood, you yourself wouldn't be willing and able to calculate correctly
WorstSkilledPlayer2026-01-13 21:01
Okay grandpa. Stop shouting at the clouds, take your meds and time to sleep 👴.
Ecchify2026-01-13 21:04
zip it up when you're done
DarkAlatreon2026-01-13 21:02
I use AI a lot, even at work. It can be very useful when properly used for things it's good at.
What I am against is content that boils down to "look guys, I asked AI a question and it gave me an answer!"
It lacks effort, accuracy and entertainment.
Ricksaw262026-01-13 18:51
You have probably better chances of becoming goku tomorrow.
Overall_Historian_802026-01-13 19:00
So you're telling me there's a chance
Hydrophobic_Stapler2026-01-13 19:17
50/50, it either happens or ~~it doesn’t~~ you get Qiqi
Overall_Historian_802026-01-13 19:19
50% of the time it works every time
Kaymish_2026-01-13 22:01
For me it seems to be Jean. Hoyo doesn't want me to get a C6 character, so they swapped me out from Qiqi once I got her to C5. I wonder who I will get once Jean is at C5.
Yuki-jou2026-01-14 02:54
Maybe you’ll join me and my constant stream of Monas.
Ashlinidor2026-01-13 19:04
Surely!
TemporaryPool61502026-01-13 19:12
Just start doing EVERY rituals you can imagine 😭
lostiming2026-01-13 19:14
Tbf, when you are lucky enough, you only need 7 pulls to c6.
DazzlingConflict90982026-01-13 19:14
The more you try and fail the more the chances of succeeding at the next try rises up
_LFKrebs_2026-01-13 19:15
If you never did that one Inazuma quest that clears the skies from Yashiori island (I think that’s the name), it’s very easy to get struck by lightning 65 million times in a row there, one C6 Columbina please
IncidentOdd56942026-01-13 19:16
IT DIDNT SAY ZERO! THE ODDS ARE STILL THERE
casper_072026-01-13 19:21
U just need 2 pull to get her C6. First pull to get C0 bina and that will not have glitter, second pull will be C1 bina with 10 glitter and you’ll convert it to fate. By the time you’re done with your seventh pull. You’ll have C6 bina and an extra 5 pulls, putting u at 7 pulls
Just need a bit of luck
Cocoatrice2026-01-13 19:26
inb4: you get C6 in 10 pull + lost to standard character afterwards
Zeus in the sky: :)
Landfall246012026-01-13 19:32
I was struck by lighting 66 million times in a row so it can happen.
True story btw.
nntnot2026-01-16 12:49
This is true I was the lighting strike
Foxintoxx2026-01-13 19:34
It says you have a chance right there .
Medyanka2026-01-13 19:36
c6 columbina is atleast 7 (!) featured copies
Negative-Slip52162026-01-13 19:40
I mean It's not Zero
Ecchify2026-01-13 19:40
ai em dashes huh, im surprised ais are still getting simple calculations wrong and making stuff up (last line) in 2026
just wasting natural resources for the love of the game ✌️
armyofonetaco2026-01-13 21:38
Bro your commenting on reddit keeping their AI alive.
Just wasting natural resources for the love of shitposting.
Neolife2026-01-13 22:54
LLMs will never be consistently good at calculations without a tool call. It's not what they're meant to do. They don't view numerical inputs the way that a calculator or program does.
imnotthinkinghard2026-01-13 19:40
There's a higher chance of you becoming a millionaire
ban9132026-01-13 19:48
All you need is a 10 pull and a dream
RenShimizu2026-01-13 19:49
But are they really complete without their fully refined 5\* weapon?
Stellin692026-01-13 20:04
Imagine being the guy who's struck over 30 million of times in a row by a thunder and saying "yeah, guess I'm halfway done"
UrGettingcooked09242026-01-13 20:09
This is useless 😂🤣😂 too much cope
RustyOrangeDog2026-01-13 20:13
Rooting for you.
PoPo5732026-01-13 20:35
I believe in you!
Dardrol72026-01-13 20:41
There's always a chance!
JJhnz122026-01-13 20:44
Just so you know the best run of reds on an actual roulette wheel had around the probilty 1.1e10 so yeah not good odds
YeetaIta2026-01-13 21:23
He says as he gets dragged away in a strait jacket.
I'm getting her C2 in one ten pull!! And then maybe go for her weapon!
A_very_smol_Lugia2026-01-14 00:38
Why are you asking ai of all things for math
seexo2026-01-14 00:47
[Chances are 50/50](https://www.youtube.com/watch?v=kt_i6DTnEok)
cartercr2026-01-14 01:12
Their math is incorrect though. You would need 7 copies, not just 6!
Diligent-Ad26412026-01-14 01:12
I mean if u believe enough anything can be achievable?
Otherwise_Praline8192026-01-14 02:26
At least 7 copies not 6.
Seangel-zero2026-01-14 02:38
You don't need to be that lucky to get C6R5.
you only need to be rich..
to the point spending money comparable to 3 units of the latest iPhone on something digital does not concern you anymore
bananaroll_2026-01-14 03:01
actually its probably even lower since you need 7 copies to c6
BigNekoTiddy2026-01-14 05:30
But it isn't zero...
Captain_Jackson2026-01-14 06:59
Excuse me, I've done inazuma quests. I'm fully aware the odds of getting struck by lightning 65 million times in a row is 100% when you're stuck in a cutscene.
DanuPellu2026-01-14 08:46
« There’s still a number » moment xD
ChaosKinZ2026-01-14 18:46
I only got struck by thunder 63,233 million times in a row one day. I was so close
HelloLellow2026-01-14 20:59
I had 2 pity and had to win my 50/50. I did a ten pull and got two Columbinas back to back.
donttrustmewithagun2026-01-18 02:51
You can tell this guy's a massive loser because his GPT is using emojis
Commentaires
P(X>=7) = P(7) + P(8) + P(9) + P(10)
Edit: Wait a second, i calculated wrong! This (3.3×10^-14) astronomically low chance is for any 5 stars, not for limited. I didnt include 50/50 and guarantee. For c6 limited char chance is 0,000000000000000257%, or 2.57×10^-16
(significantly more effort but whatever)
oh and also capturing radiance which lowers the chance of 3 consecutive 50-50 losses, but the effect of any of these is minimal
your 0.994^3 or 0.997^3 assumes that the others can't be 5 stars, which they can, so it should just be 1^3 which is 1 (oh right you did say it earlier)
P(7) chances to c6 is 0.78%(1/128 i wrote earlier, need 7 wins), P(8) chances to c6 is 6.25%(1/16 cause 2×0.5\^8 of 8 wins and 7 wins and 1 lose at the end, and 7 variants of 0.5\^7 with guarantee in the middle, with sum of 1/256+1/256+7/128=16/256=1/16 or 6.25%), p(9) and p(10) contribution is negligible so i just ignored it. Just sum it and real final chance to c6 in 10 pulls is (3.3×10\^-14)×0.78% + (7.5×10\^-17)×6.25% = 2.625×10\^-16
edit: recalculated, firstly got wrong result. Btw, why this time formatting doesn't work??? Btw my brain is fried, i edited it like 10 ten time after each recalculation.
[This](https://docs.google.com/spreadsheets/d/1XXbShUXR-RLmDxTZmXzl_Z7ye5ve1lfaVmLHXsKFvxg/edit?usp=sharing) was my calculation, let me know if I did this wrong. Technically my calculation doesn't take into account the probability of getting a guaranteed after losing 5050 on one of the 3 pulls that we can spare, but I don't think that changes the results in a meaningful way.
I'm not sure how the pity system comes into play, we are always before any kind of soft pity, so every pull should have a chance of 0.3% for the 5 star we want.
I see you used binomial distribution which is also right thing, but may underestimate chances in compare to bernoulli(not this time). You just need to use right chances and results become similar with mine. (i made copy of your doc, changed chance from 0.3% to 0.6%, included pity and got 2.625e-16)
edit: forget what i wrote in edit earlier, right calculations is above
Code execution result
Five Stars | Binomial Prob | Chance of 7+ Event chars | Contribution
--------------------------------------------------------------------------------
7 | 3.29913e-14 | 0.00781 | 2.57744e-16
8 | 7.46784e-17 | 0.06250 | 4.66740e-18
9 | 1.00172e-19 | 0.22656 | 2.26953e-20
10 | 6.04662e-23 | 0.50000 | 3.02331e-23
--------------------------------------------------------------------------------
Total Probability: 2.6243449537348796e-16
One in 3,810,474,680,841,151
the number it gave is wrong, literally just input that on a calculator to check. and the last statement about lightning odds don't make any sense either lmao
cmon man, let's use our brains a little. literally no one besides billionaires benefit from dumb mindsets like yours
and 10 pull is an arbitrary number anyway, if you want a simple calculation that doesn't need any knowledge about maths just calculate odds for 7 in a row instead.
ultimately it doesn't matter either way since the answer is some super small number that in practical terms is just 0, but if someone actually wanted to know the real answer to that question for some reason, it's not a straightforward calculation
edit: at least 7, not exactly 7
What I am against is content that boils down to "look guys, I asked AI a question and it gave me an answer!"
It lacks effort, accuracy and entertainment.
Just need a bit of luck
Zeus in the sky: :)
True story btw.
just wasting natural resources for the love of the game ✌️
Just wasting natural resources for the love of shitposting.
C0 -> C1 -> C2 -> C3 -> C4 -> C5 -> C6
you only need to be rich..
to the point spending money comparable to 3 units of the latest iPhone on something digital does not concern you anymore